The centers of two masses are 5m apart. The masses exert a gravitational force of

2.4x10^-4 N on each other. Find the masses if one is three times the mass of the other.


1
Expert's answer
2021-10-24T18:19:45-0400

Given:

"F=2.4*10^{-4}\:\rm N"

"m_1=3m_2"

"r=5\:\rm m"


The gravity force is given by:


"F=G\frac{m_1m_2}{r^2}"

We get

"F=G\frac{3m_2^2}{r^2}"

"m_2=\sqrt{\frac{r^2F}{3G}}=\sqrt{\frac{5^2*2.4*10^{-4}}{3*6.67*10^{-11}}}=5476\:\rm kg"

"m_1=3m_2=3*5476\:\rm kg=16428\:\rm kg"